Accelerator Update for July 30, 2008 - August 1, 2008
Wednesday July 30
The day shift began with Operations monitoring the Tevatron (TeV) store 6324 (56.2E30) and the Recycler (RR) stash (91.9E10), with the Antiproton Source off due to D:H717 repairs, with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the SciBooNE / MiniBooNE (BNB), NuMI, and Meson MTest experiments taking beam.
Operations disabled the MTest beam at 8:50 AM.
At 9:02 AM, Operations halted all beam. Operators began preparing the Pbar transport line and rings for access. Interlock technicians will test the Booster Critical Device Controller (CDC) before any keys will be issued.
Operations reported at 9:08 AM that Booster beam was set up to go to the absorber.
Interlock techs reported at 10:13 AM that they were testing MI-31 and the Pelletron.
The Pbar access began at 10:26 AM.
Pbar experts reported at 10:27 AM that H717 was electrically sound.
Operations prepared NuMI for access at 10:32 AM. NuMI was out of access an hour later.
Pbar experts reported at 1:14 PM that they were beginning maintenance on the Pbar lithium lens and cooling systems.
Pbar came out of access at 4:34 PM.
At 4:38 PM, a Linac RF station (LRF4) tripped off and would not reset. Experts investigated.
LCW experts reported at 4:41 PM that LCW was flowing through Pbar.
A Linac expert reported that he found a bad 24-volt power supply for LRF4. He had the station back on line by 6:05 PM.
Operations established beam to NuMI and BNB at 7:01 PM.
Operators prepared Pbar for access at 7:38 PM. H717 alarmed on magnet over temperature and water flow fault alarms. An expert investigated.
Pbar came out of access at 8:02 PM. Power supply experts were checking out the power supply.
Operations established beam to NuMI and BNB at 8:03 PM.
Operations reported at 8:51 PM that H717 was operating normally. The expert reported that he found a bad opto-coupler in the power supply.
Pbar resumed stacking at 8:59 PM.
Operations established beam to MTest around 10 PM.
At 10:08 PM, LRF4 tripped off. It reset.
Operations reported at 11:57 PM that the Recycler had stashed once during the evening shift.
Thursday July 31
The midnight shift began with Operations monitoring store 6324 (30.17E30) and the stash (121.5E10), with Pbar stacking (36.9E10), with the BNB, NuMI, and MTest experiments taking beam.
At 12:37 AM, the vacuum pump for LRF3 tripped off. It reset.
At 3:09 AM, a Booster vacuum pump (IPRF17) tripped off and caused BRF7 & 8 to trip off. Everything reset.
Operations terminated store 6324 at 4:11 AM.
Operations put the TeV into shot setup at 4:17 AM.
At 4:58 AM, MIRF14 tripped off. It reset.
Operations established store 6325 at 6:53 AM with an initial luminosity of 182.1E30.
At 7:56 AM, LRF2 tripped off. It reset.
Operations reported at 7:59 AM that the Recycler had stashed four times during the midnight shift.
At 9:01 AM, Operations prepared MI-31 for an access. An Electron Cooling expert reported that he needed to increase the motor current for the SF6 blower.
Operations reported at 9:06 AM that FESS personnel were beginning work on the MI-10 air conditioner.
MI-31 came out of access at 9:33 AM.
At 9:51 AM, Operations reported that four Pbar Debuncher vacuum pumps (IP307, IPPMA, IPPMB, and IP902) tripped off and would not reset. Operators investigated.
Operations disabled the MTest beam at 9:58 AM.
A RF expert reported at 10:41 AM that FESS had reconfigured the pond pumps at MI-60.
At 1:38 PM, Operations turned off the BNB beam line due to high MI LCW temperatures.
Operations resumed sending beam to BNB at 2:52 PM.
Operations reported at 3:27 PM that the MI-20 LCW pumps were realigned and ready to be turned on.
Operations reported at 3:58 PM that the Recycler had stashed six times during the day shift.
At 6:59 PM, the MI Power Supply interlock Loop tripped off. A VCB had opened. Everything reset.
At 8:12 PM, Operations turned off the BNB line to lower the MI LCW temperature.
Operations established beam to MTest at 8:29 PM.
At 8:41 PM, Operations disabled the MTest beam due to the rise in MI LCW temperatures.
Operations terminated store 6325 at 10:02 PM.
Operations put the TeV into shot setup at 10:13 PM.
At 11:02 PM, Operations reported an Electron Cooling energy error that was slowing the antiproton cooling process.
At 11:27 PM, Operations reported a problem with a Pbar Debuncher RF station (DRF1-6). It was tripping off on a driver fault and then resetting itself. This occurred repeatedly over a 10-minute period and then stabilized.
Operations turned on the MI-20 LCW pumps at 11:37 PM.
At 11:53 PM, LRF4 tripped off. It reset.
Operations reported at 11:56 PM that the Recycler had stashed four times during the evening shift.
Friday August 1
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (92.9E10), with Pbar stacking (57.5E10), and with the BNB, NuMI, and MTest experiments taking beam.
Operations established store 6326 at 12:01 AM with an initial luminosity of 277.56E30.
Operations established beam to MTest at 12:03 AM.
At 2:17 AM, a Linac Klystron RF station (KRF3) tripped off on a water flow indication and would not reset. Operations contacted a Linac expert.
Operators brought KRF3 back online at 3:22 AM.
At 7:18 AM, LRF4 tripped off. It reset.
Operations reported at 8 AM that the Recycler had stashed four times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today and for the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 08/01/2008 email Fermilab|
|Security, Privacy, Legal|